# Noise-augmented Chaotic Ising Machines for Combinatorial Optimization and Sampling

### v1

@@ Abstract
The rise of domain-specific computing has led to great interest in Ising machines, dedicated hardware accelerators tailored to solve combinatorial optimization and probabilistic sampling problems. A key element of Ising machines is stochasticity, which enables a wide exploration of configurations, thereby helping avoid local minima. Here, we evaluate and improve the previously proposed concept of coupled chaotic bits (c-bits) that operate without any explicit stochasticity. We show that augmenting chaotic bits with stochasticity leads to better algorithmic scaling in combinatorial optimization problems, comparable to the performance of probabilistic bits (p-bits) which have explicit randomness in their update rules. We first demonstrate that c-bits surprisingly follow the quantum Boltzmann law in a 1D transverse field Ising model, despite the lack of explicit randomness. We then show that c-bits exhibit critical dynamics similar to those of stochastic p-bits in 2D Ising and 3D spin glass models, with promising potential to solve challenging optimization problems. Finally, we propose a noise-augmented version of coupled c-bits via the powerful adaptive parallel tempering algorithm (APT). The noise-augmented c-bit algorithm outperforms fully deterministic c-bits running versions of the simulated annealing algorithm. Chaotic Ising machines closely resemble coupled oscillator-based Ising machines, as both schemes exploit nonlinear dynamics for computation. Oscillator-based Ising machines may greatly benefit from our proposed algorithm, which runs replicas at constant temperature, eliminating the need to globally modulate coupling strengths. Mixing stochasticity with deterministic c-bits creates a powerful hybrid computing scheme that can bring benefits in scaled, asynchronous, and massively parallel hardware implementations.
@@ I.   Introduction
As Moore’s law stagnates, there is growing interest in unconventional computing schemes and hardware accelerators. Among these emerging technologies are Ising Machines [1], employing devices such as probabilistic bits (p-bits) [2, 3, 4, 5], coupled oscillators [6, 7, 8, 9, 10, 11, 12], photonic devices [13, 14, 15, 16], and other nonlinear elements [17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. Ising machines have traditionally focused on providing scaling and prefactor improvements in solving combinatorial optimization problems [27, 28, 29]. Efforts to link Ising machines to generative machine learning [30, 31, 32, 33, 34] and quantum simulation problems are also underway [35, 36, 37]. Here, we evaluate and improve an earlier Ising machine concept based on chaotic bits (c-bits), originally proposed in Ref. [38]. Chaotic bits resemble coupled oscillators, following a set of deterministic neuron update rules without explicit randomness. However, unlike oscillator-based Ising Machines that encode the Ising spin in the continuous phase of oscillators, c-bits make use of oscillating billiard balls that set a latch to a $+1$ or $-1$ state. As such, chaotic bits do not require cumbersome subharmonic injection locking schemes to binarize naturally continuous phase variables.
The potential hardware implementation of chaotic Ising machines without any explicit random number generators, while retaining similarity to the mathematics of p-bits, is appealing. However, as we demonstrate in this paper, c-bits without any randomness cannot straightforwardly employ the most powerful Monte Carlo algorithms, such as Adaptive Parallel Tempering, which offer better algorithmic scaling for the Circuit SAT and 3D Spin Glass problems. We propose and evaluate a hybrid computing scheme wherein deterministic c-bits are augmented with stochasticity for improved performance in combinatorial optimization and sampling. The approach we propose may be applied to similar devices, such as coupled oscillators, where a continuous global modulation of analog coupling elements may be inconvenient.
@@
where $m_{i}\in\{-1,+1\}$ and $\text{rand}_{u}(-1,1)$ is a random uniform distribution between $-1$ and $1$ (FIG. 1a, c). $\beta$ is inverse temperature. $J_{ij}$ and $h_{i}$ represent weights and biases. At equilibrium, p-bit networks sample from the Boltzmann distribution, given by:
@@
where $\{m\}$ is a spin configuration, $Z$ is the partition function, and $E(\{m\})$ is the energy of a spin configuration. Chaotic bits can be formulated to follow similar equations as p-bits, that describe an oscillating billiard ball and a latch:
@@
where $x_{i}$ represents the position of a billiard between the boundaries $x_{i}=-1$ and $x_{i}=+1$. The billiard has a tunable slope, ${dx_{i}}/{dt}$, which is dependent on the input $I_{i}$. When the billiard $x_{i}$ reaches $-1$ or $+1$, the c-bit state $m_{i}$ is latched to that value, and the billiard changes direction (FIG. 1b). Note an interesting parallel to p-bits. In Gibbs sampling, simultaneously updating connected nodes leads to pathological oscillations [39, 40]. For c-bits, the differential equation formulation in continuous time prevents simultaneous updates naturally, as long as the time difference between subsequent billiard ball arrivals is greater than the time it takes to update a latch. When c-bits are discretized in software, latches are set instantaneously within a single time step, $\Delta t$, chosen to solve the differential equations. In Section VII, we briefly discuss implications of this update mechanism for eventual hardware implementations. It is important to note that p-bits and c-bits have different notions of time. There is no clear correspondence between a p-bit Monte Carlo sweep (MCS), in which Eqs. 1 and 2 are sequentially solved for each p-bit in the network, and the time variable $t$ of a c-bit network. In this study, we sample the state $\{m\}$ of a continuously updating c-bit network at natural numbers $t=\{0,1,2,\ldots,t_{a}\}$.
Individual p-bits in a system observe the following relationship:
@@
Intuitively, c-bits try to achieve the same by modifying the lifetime of up and down trajectories of the billiard balls, conditioned on the state of their neighbors (FIG. 1b):
@@
where $\tau$ is the lifetime of a c-bit at its current state. The key point is that $\tau$ are deterministic, unlike a p-bit whose state is stochastically sampled at each update. Eq. 5 is not the only valid c-bit definition [41]. As long as the ratio in Eq. 6 is observed, a c-bit could be defined using different velocities.
@@
This intuitive explanation does not show that c-bit networks sample from the Boltzmann distribution as stochastic p-bits do. Theoretical work has shown convergence to the Boltzmann distribution for simple toy models [42]. Experimental work suggests that c-bit networks seem to closely approximate the Boltzmann distribution for 2D Ising models of small sizes [41]. Furthermore, it has been shown that c-bits match stochastic p-bits for equilibrium statistics on the 2D ferromagnetic Ising model [38] and the Potts model [41]. c-bits have been used to determine the critical temperature ($T_{c}$) and critical exponents ($\nu$, $\beta$, and $\gamma$) of the Ising universality class [41]. Simulated annealing has been demonstrated with c-bits [38, 43]. Additionally, CMOS and analog VLSI implementations have been proposed [44, 45]. Nonetheless, the theoretical equivalence of c-bits to p-bits without explicit noise is far from clear. Moreover, the initial phase randomization and rounding errors in the solution of ODEs (or physical noise in real systems) might be responsible for this striking correspondence.
@@
Here, we subject c-bits to more stringent tests on three important problem classes, comparing their performance to p-bits. These tests involve a quantum problem, criticality in 3D spin-glasses, and hard combinatorial optimization in the form of integer factorization with invertible logic gates. We conclude that while c-bits may perform similarly to p-bits in algorithms such as simulated annealing, they may benefit from added stochasticity, either through their update rules (for example, by adding jitter to the billiard periods), or by adding explicit random swaps used in powerful tempering algorithms.
@@
Most existing work employs a definition of ${dx_{i}}/{dt}$ that grows exponentially at cold temperatures ($\beta\gg 1$), with the exception of one previous study [41]. This is problematic because optimization algorithms such as simulated annealing and adaptive parallel tempering require cold temperatures which may result in unbounded slopes. In this study, we employ a c-bit definition that has a maximum billiard speed of 2 ($\sum\tau(m_{i}=\pm 1)=2$, see, Eq. 5), suitable for software and hardware implementation. Unless otherwise specified, we employ the Euler’s Method with a step of $\Delta t=0.1$ to simulate the dynamics of coupled c-bits.
@@
Finally, in all of our experiments, the initial phases of the billiard balls and the states of the latches are randomized. The asynchronous dynamics of the billiard balls, coupled with this phase randomization ensures nonsimultaneous updates with quasi-random arrivals. In addition, rounding errors and the choice of a naive integration scheme provide additional noise in our simulations. We believe these seem to be the key reasons behind the near-equivalence of c-bits to p-bits where explicit and implicit inclusion of noise seems to play a key role.
@@ II.   Sampling from the 1D Transverse Field Ising Hamiltonian with c-bits
where $\sigma_{i}^{x}$ and $\sigma_{i}^{z}$ are Pauli spin matrices at site $i$. The interaction weight $J_{ij}=+2$ for nearest neighbors (with a periodic boundary condition) and $J_{ij}=0$ otherwise. $h_{i}=+1$, representing a symmetry breaking magnetic field in the $+\hat{z}$ direction. The Suzuki-Trotter transformation, [47, 35] maps the 1D quantum Hamiltonian to the following 2D classical Hamiltonian:
@@
where $J_{\perp}=-{1}/{2\beta}\ln{\tanh{\left({\beta\Gamma_{x}}/{R}\right)}}$. Note the unusual dependence of this coupling to the inverse temperature, $\beta$.
@@ III.   Critical Scaling Dynamics: c-bits vs p-bits
Fig. 3: (a) Residual energy density at the critical point ($\rho_{E}^{c}$) as a function of annealing time for a $50\times 50$ ferromagnetic Ising model lattice with nonperiodic boundary conditions ($J_{ij}=+1$). We perform linear regression on the 5 leftmost points of each data series to avoid finite-size effects. Error bars show $95\%$ bootstrap confidence intervals. (b) $\rho_{E}^{c}$ vs. annealing time for cubic spin-glass instances ($L=11$, $J_{ij}\in\{-1,+1\}$). (c) $\rho_{E}^{c}$ as a function of inverse annealing velocity for the 2D ferromagnetic Ising model, using p-bits. Different annealing schedules are shown in a data collapse by considering the quench rate at the critical point, $\alpha$. (d) $\rho_{E}^{c}$ vs. inverse annealing velocity for the 2D ferromagnetic Ising model, using c-bits.
@@
Although there is no obvious correspondence between a Monte Carlo sweep for behavioral models of p-bits and the time step of a c-bit, power-law exponents can be compared objectively since they do not depend on prefactors. FIG. 3a shows that c-bits and p-bits demonstrate similar critical scaling dynamics on the 2D ferromagnetic Ising model. Linear regression is performed on the 5 leftmost points in order to avoid finite-size effects. Both p-bit and c-bit curves exhibit downward concavity, but different time scales mean that this concavity does not manifest at the same time $t_{a}$ for p-bits and c-bits. We provide the results of using Euler’s method in steps of $\Delta t=0.1$ and $\Delta t=0.01$. The c-bit curve using $\Delta t=0.01$ appears to follow a similar power law scaling, but with a steeper slope $|\kappa_{c}|$. The reason for this discrepancy is not clear. On the one hand, smaller $\Delta t$ implies higher theoretical precision, while on the other hand, smaller $\Delta t$ leads to more round-off errors due to double precision. It must also be noted that a lower $\Delta t$ requires more computational effort, and in hardware, $\Delta t$ may have a lower bound dictated by the physical latch required to set the state of the oscillator.
@@
FIG. 3b shows a similar experiment conducted on select instances of the 3D spin glass problem, studied by D-Wave [50]. We consider 300 different spin glass instances on a cubic lattice with side length $L=11$ and $J_{ij}\in\{-1,+1\}$. Analogous to FIG. 3a, for a given annealing time $t_{a}$, we geometrically increase $\beta$ to the inverse critical temperature $1/T_{c}$. For 3D spin glasses, $T_{c}$ is approximately calculated as $\approx 1.1$ [53]). For each of 300 problem instances, we take an ensemble average over 10 random seeds. We again observe similar power-law scaling. Our p-bit results match the numerical results in the literature, with $\kappa_{c}\approx 0.51$ [50]. However, the values of $\kappa_{c}$ do not seem to be similar when comparing between p-bits and c-bits. The $\Delta t$ dependence of c-bit slopes also is an indication that the underlying dynamics for c-bits might depend on separating near arrivals that latch a state (where Eq. 2 is always assumed to be infinitely fast). We conclude that while c-bits seem to obey a similar power law as p-bits, their dynamics seems different in subtle ways.

### v2

@@ Abstract
Ising machines, hardware accelerators for combinatorial optimization and probabilistic sampling problems, have gained significant interest recently. A key element is stochasticity, which enables a wide exploration of configurations, thereby helping avoid local minima. Here, we refine the previously proposed concept of coupled chaotic bits (c-bits) that operate without explicit stochasticity. We show that augmenting chaotic bits with stochasticity enhances performance in combinatorial optimization, achieving algorithmic scaling comparable to probabilistic bits (p-bits). We first demonstrate that c-bits follow the quantum Boltzmann law in a 1D transverse field Ising model. We then show that c-bits exhibit critical dynamics similar to stochastic p-bits in 2D Ising and 3D spin glass models, with promising potential to solve challenging optimization problems. Finally, we propose a noise-augmented version of coupled c-bits via the adaptive parallel tempering algorithm (APT). Our noise-augmented c-bit algorithm outperforms fully deterministic c-bits running versions of the simulated annealing algorithm. Other analog Ising machines with coupled oscillators could draw inspiration from the proposed algorithm. Running replicas at constant temperature eliminates the need for global modulation of coupling strengths. Mixing stochasticity with deterministic c-bits creates a powerful hybrid computing scheme that can bring benefits in scaled, asynchronous, and massively parallel hardware implementations.
@@ I.   Introduction
As Moore’s law stagnates, there is growing interest in unconventional computing schemes and hardware accelerators. Among these emerging technologies are Ising machines [1], employing devices such as probabilistic bits (p-bits) [2, 3, 4, 5], coupled oscillators [6, 7, 8, 9, 10, 11, 12], photonic devices [13, 14, 15, 16], and other nonlinear elements [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Ising machines have traditionally focused on providing scaling and prefactor improvements in solving combinatorial optimization problems [29, 30, 31]. Efforts to link Ising machines to generative machine learning [32, 33, 34, 35, 36] and quantum simulation problems are also underway [37, 38, 39]. Here, we evaluate and improve an earlier Ising machine concept based on chaotic bits (c-bits), originally proposed in Ref. [40]. Chaotic bits resemble coupled oscillators, following a set of deterministic neuron update rules without explicit randomness. However, unlike oscillator-based Ising Machines that encode the Ising spin in the continuous phase of oscillators, c-bits make use of oscillating billiard balls that set a latch to a $+1$ or $-1$ state. As such, chaotic bits do not require cumbersome subharmonic injection locking schemes to binarize naturally continuous phase variables.
The potential hardware implementation of chaotic Ising machines without any explicit random number generators, while retaining similarity to the mathematics of p-bits, is appealing. However, as we demonstrate in this paper, c-bits without any randomness cannot straightforwardly employ the most powerful Monte Carlo algorithms, such as Adaptive Parallel Tempering, which offer better algorithmic scaling for the Circuit SAT and 3D Spin Glass problems. We propose and evaluate a hybrid computing scheme wherein deterministic c-bits are augmented with stochasticity for improved performance in combinatorial optimization and sampling. The approach we propose may be applied to similar devices, such as certain classes of coupled oscillators, where a continuous global modulation of analog coupling elements may be inconvenient.
@@
[…]
@@
where $\{m\}$ is a spin configuration, $Z$ is the partition function, and $E(\{m\})$ is the energy of a spin configuration. Chaotic bits follow deterministic update rules, inspired by p-bit equations, which model an oscillating billiard ball and a latch. We reformulate the original chaotic bit equations [40, 41] to emphasize their similarity to p-bits:
@@
where $x_{i}$ represents the position of a billiard between the boundaries $x_{i}=-1$ and $x_{i}=+1$. The billiard has a tunable slope, ${dx_{i}}/{dt}$, which is dependent on the input $I_{i}$. When the billiard $x_{i}$ reaches $-1$ or $+1$, the c-bit state $m_{i}$ is latched to that value, and the billiard changes direction (FIG. 1b). Note that in numerical or hardware implementations of c-bits, the billiard state $x_{i}$ in Eq. 5 can slightly exceed +1 or --1 due to discretization errors.
The c-bit definition can be motivated from the Boltzmann distribution, analogous to p-bits and Gibbs sampling. Consider a single spin $m_{i}$ with a fixed set of neighbors. For p-bits, combining Eq. 2 and Eq. 3, the conditional probability of $m_{i}$ being ±$\pm 1$ is expressed as:
$p(m_{i}=+1)&=\frac{1}{Z}\exp{[\beta(I_{i}-E_{\text{rest}})]}\\ p(m_{i}=-1)&=\frac{1}{Z}\exp{[\beta(-I_{i}-E_{\text{rest}})]}\end{aligned}$ (6)
where $Z$ is the partition function and $E_{\text{rest}}$ is the part of the global energy (or the environment) that is not dependent on p-bit $i$. When expressed as a ratio of probabilities, Eq. 6 becomes:
@@
which describes the behavior of a single p-bit given fixed neighbors.
As a deterministic unit, c-bit updates do not involve probabilities, but rather they try to achieve the same sampling ratio of $+1$’s to $-1$’s by modifying the lifetime of up and down trajectories of the billiard ball, conditioned on the state of its neighbors (FIG. 1b):
@@
where $\tau$ is the lifetime of a c-bit at its current state ($m_{i}=\pm 1$). In other words, c-bits are defined such that the time average of the c-bit state $\langle m_{i}\rangle=\tanh{(\beta I_{i})}$, analogous to a p-bit. The key point is that $\tau$ are deterministic, unlike a p-bit whose state is stochastically sampled at each update.
The comparison of p-bits (Eq. 1 and Eq. 2) with c-bits (Eq. 5 and Eq. 2) is fundamental. In the case of p-bits, the fundamental theorem of Markov Chains [42, 43] ensure that a network of p-bits updated sequentially via Gibbs sampling will eventually reach the Boltzmann distribution, defined by Eq. 3. Ultimately, this is what enables mapping problems of interest to p-bit networks and then solving them with algorithms such as simulated annealing and parallel tempering. Networks of c-bits cannot be proved to sample exactly from the Boltzmann distribution due the lack of probabilistic and serial updates.
Fig. 2: (a) Distribution of spin configurations for a 5 p-bit full adder, obtained analytically from the Boltzmann distribution. (b) Analytical distribution for p-bits employing parallel updates, as opposed to sequential Gibbs sampling [5]. (c) Numerical results for parallel p-bit updates, collected over $10^{5}$ Monte Carlo sweeps. (d) c-bits exhibit similar pathological behavior when their phases $x_{i}$ and states $m_{i}$ are all initialized to the same value, $+1$. Samples are collected over $10^{5}$ time steps.
Note the following parallel to p-bits. It is well-known that in Gibbs sampling, simultaneously updating connected nodes leads to pathological oscillations [44, 45], preventing easy parallelization of the chain. For c-bits, it seems that the differential equation formulation in continuous time prevents simultaneous updates naturally, as long as the time difference between subsequent billiard ball arrivals is greater than the time it takes to update a latch. When c-bits are discretized in software, latches are set instantaneously within a single time step, $\Delta t$, chosen to solve the differential equations.
To better understand how critical serial updating is for c-bits, we construct a simple example, with 5 c-bits (or 5 p-bits), interconnected such that their low energy states represent the truth table of a full adder, given by the interactions [46, 47]:
$J_{FA}=\begin{bmatrix}0&-1&-1&1&2\\ -1&0&-1&1&2\\ -1&-1&0&1&2\\ 1&1&1&0&-2\\ 2&2&2&-2&0\end{bmatrix}$
with no bias terms, $h=0$. The columns correspond to carry in, A, B, sum and carry out bits, respectively. FIG. 2a shows the analytically obtained Boltzmann distribution from Eq. 3 at $\beta=0.5$, which can be numerically approximated by sequential p-bit updates.
FIG. 2b shows the probability distribution of 32 possible spin configurations in a 5 p-bit full adder when p-bits are updated in parallel rather than sequentially. The steady-state distribution corresponding to this Markov chain can be obtained analytically [5, 43] and describes the numerical sampling (shown in FIG. 2c) well.
Remarkably, FIG. 2d shows that similar pathological oscillations between states 3 and 28 are observed in c-bit networks when the initial states are synchronized by making all $x_{i}$ and $m_{i}$ equal in the beginning. These oscillations go away if the initial phases are sufficiently separate and the integration step, $\Delta$t is small. We conclude that c-bit networks implicitly rely on sequential updating for correct operation, despite their continuous time and massively parallel dynamics. While simultaneous c-bit updates in an integration step $\Delta t$ may be considered an artifact of software discretization, hardware implementations will have similar restrictions due to finite setup/hold times and synapse delays associated with calculating $I_{i}$. In Section VII, we briefly discuss implications of this update mechanism for eventual hardware implementations.
It is important to note that p-bits and c-bits have different notions of time. There is no clear correspondence between a p-bit Monte Carlo sweep (MCS), in which Eqs. 1 and 2 are sequentially solved for each p-bit in the network, and the time variable $t$ of a c-bit network. In this study, we sample the state $\{m\}$ of a continuously updating c-bit network at natural numbers $t=\{0,1,2,\ldots,t_{a}\}$. Because of these different notions of time, in order to draw comparisons between p-bit and c-bit algorithms, we shall employ power law and exponential scaling arguments, which hold independent of prefactors. Furthermore, we apply similar algorithmic parameters to both p-bits and c-bits to ensure fair comparisons.
@@
Theoretical work has shown convergence to the Boltzmann distribution for simple toy models of 2 c-bits [48]. However, it has not been determined that larger c-bit networks necessarily sample from the Boltzmann distribution as stochastic p-bits do. Experimental work suggests that c-bit networks seem to closely approximate the Boltzmann distribution for 2D Ising models of small sizes [41]. Furthermore, it has been shown that c-bits match stochastic p-bits for equilibrium statistics on the 2D ferromagnetic Ising model [40] and the Potts model [41]. c-bits have been used to determine the critical temperature ($T_{c}$) and critical exponents ($\nu$, $\beta$, and $\gamma$) of the Ising universality class [41]. Simulated annealing has been demonstrated with c-bits [40, 49]. Additionally, CMOS and analog VLSI implementations have been proposed [50, 51]. Nonetheless, the theoretical equivalence of p-bits to c-bits without explicit noise is far from clear. Moreover, the initial phase randomization and rounding errors in the solution of ODEs (or physical noise in real systems) might be responsible for this striking correspondence.
@@
Here, we subject c-bits to more stringent tests on three important problem classes, comparing their performance to p-bits. These tests involve a quantum problem, criticality in 3D spin-glasses, and hard combinatorial optimization in the form of integer factorization with invertible logic gates. We conclude that while c-bits perform similarly to p-bits in algorithms such as simulated annealing, c-bits also benefit from added stochasticity, particularly through random swaps used in powerful tempering algorithms.
@@
Most existing work employs a definition of ${dx_{i}}/{dt}$ that grows exponentially at cold temperatures ($\beta\gg 1$), with the exception of one previous study [41]. This is problematic because optimization algorithms such as simulated annealing and adaptive parallel tempering require cold temperatures which may result in unbounded slopes. In this study, we employ a c-bit definition that has a maximum billiard speed of 2 (see Eq. 5), suitable for software and hardware implementation. Unless otherwise specified, we employ the Euler’s Method with a step of $\Delta t=0.1$ to simulate the dynamics of coupled c-bits.
@@
Finally, in all of our experiments, the initial phases of the billiard balls and the states of the latches are randomized. The asynchronous dynamics of the billiard balls, coupled with this phase randomization ensures nonsimultaneous updates with quasi-random arrivals. In addition, rounding errors and the choice of a naive integration scheme provide additional noise in our simulations. We believe these seem to be the key reasons behind the near-equivalence of c-bits to p-bits, where explicit and implicit inclusion of noise seems to play a central role.
@@ II.   Sampling from the 1D Transverse Field Ising Hamiltonian with c-bits
where $\sigma_{i}^{x}$ and $\sigma_{i}^{z}$ are Pauli spin matrices at site $i$. Note that each of these is $2^{N}\times 2^{N}$ matrix and is computed as $\sigma_{i}^{\alpha}=I_{2}^{\otimes i-1}\otimes\sigma^{\alpha}\otimes I_{2}^{\otimes N-i}=I_{2}\otimes\ldots\otimes\sigma^{\alpha}\otimes\ldots\otimes I_{2}$ with $\alpha\in\{x,z\}$ where $\otimes$ denotes the ‘Kronecker product’, $I_{2}$ is the $2\times 2$ identity matrix, $\sigma^{z}$ and $\sigma^{x}$ are Pauli spin matrices (only at the $i$-th term of the product):
$\sigma^{z}=\left[\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right]\quad\text{and}\quad\sigma^{x}=\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right]$ (14)
The interaction weight $J_{ij}=+2$ for nearest neighbors (with a periodic boundary condition) and $J_{ij}=0$ otherwise. $h_{i}=+1$, representing a symmetry breaking magnetic field in the $+\hat{z}$ direction. The Suzuki-Trotter transformation [53, 37] maps the 1D quantum Hamiltonian to the following 2D classical Hamiltonian:
@@
where $J_{\perp}=-{1}/{2\beta}\ln{\tanh{\left({\beta\Gamma_{x}}/{R}\right)}}$. $R$ represents the total number of system replicas after applying the quantum-to-classical mapping. Note the unusual dependence of this coupling to the inverse temperature, $\beta$.
@@
$\sum_{i}\sigma_{i}^{z}$ in the numerator is the corresponding quantum operator for measuring spins along the z-direction. Notice that the argument of the ‘exp ($\cdot$)’ in Eq. 17 is a matrix. Evaluating this equation becomes intractable for large $N$, as it involves taking the exponential of a $2^{N}\times 2^{N}$ matrix. When $N$ is small like the system in our example, one can compute exponential of a matrix directly by explicitly writing down $H_{Q}$ and calculating its exponential. Also, to obtain numerically stable results at low temperatures (high $\beta$), we first diagonalize the Hamiltonian and subtract the ground-state energy from the diagonals which does not affect any observable quantities.
@@ III.   Critical Scaling Dynamics: c-bits vs p-bits
Fig. 4: (a) Residual energy density at the critical point ($\rho_{E}^{c}$) as a function of annealing time for a $50\times 50$ ferromagnetic Ising model lattice with nonperiodic boundary conditions ($J_{ij}=+1$). We perform linear regression on the 5 leftmost points of each data series to avoid finite-size effects. Error bars show $95\%$ bootstrap confidence intervals. The inset plot shows short-term exponential divergence of c-bit systems with slightly perturbed initial conditions. The slope represents a positive (maximum) Lyapunov exponent, implying chaotic behavior. (b) $\rho_{E}^{c}$ vs. annealing time for cubic spin-glass instances ($L=11$, $J_{ij}\in\{-1,+1\}$). (c) $\rho_{E}^{c}$ as a function of inverse annealing velocity for the 2D ferromagnetic Ising model, using p-bits. Different annealing schedules are shown in a data collapse by considering the quench rate at the critical point, $\alpha$. (d) $\rho_{E}^{c}$ vs. inverse annealing velocity for the 2D ferromagnetic Ising model, using c-bits.
@@
Although there is no obvious correspondence between a Monte Carlo sweep for behavioral models of p-bits and the time step of a c-bit, power-law exponents can be compared objectively since they do not depend on prefactors. FIG. 4a shows that c-bits and p-bits demonstrate similar critical scaling dynamics on the 2D ferromagnetic Ising model. Linear regression is performed on the 5 leftmost points in order to avoid finite-size effects. Both p-bit and c-bit curves exhibit downward concavity, but different time scales mean that this concavity does not manifest at the same time $t_{a}$ for p-bits and c-bits. We provide the results of using Euler’s method in steps of $\Delta t=0.1$ and $\Delta t=0.01$. The c-bit curve using $\Delta t=0.01$ appears to follow a similar power law scaling, but with a steeper slope $|\kappa_{c}|$. The reason for this discrepancy is not clear. On the one hand, smaller $\Delta t$ implies higher theoretical solver precision, while on the other hand, smaller $\Delta t$ leads to more round-off errors due to double precision. It must also be noted that a lower $\Delta t$ requires more computational effort, and in hardware, $\Delta t$ may have a lower bound dictated by the physical latch required to set the state of the oscillator.
@@
FIG. 4b shows a similar experiment conducted on select instances of the 3D spin glass problem, studied by D-Wave [56]. We consider 300 different spin glass instances on a cubic lattice with side length $L=11$ and $J_{ij}\in\{-1,+1\}$. Analogous to FIG. 4a, for a given annealing time $t_{a}$, we geometrically increase $\beta$ to the inverse critical temperature $1/T_{c}$. For 3D spin glasses, $T_{c}$ is approximately calculated as $\approx 1.1$ [59]). For each of 300 problem instances, we take an ensemble average over 40 random seeds. We again observe similar power-law scaling. Our p-bit results match the numerical results in the literature, with $\kappa_{c}\approx 0.51$ [56]. However, the values of $\kappa_{c}$ do not seem to be similar when comparing between p-bits and c-bits. The $\Delta t$ dependence of c-bit slopes also is an indication that the underlying dynamics for c-bits might depend on separating near arrivals that latch a state (where Eq. 2 is always assumed to be infinitely fast). We conclude that while c-bits seem to obey a similar power law as p-bits, their dynamics seem different in subtle ways.
One feature of chaotic dynamics is sensitivity to initial conditions. This can be tested by calculating Lyapunov exponents which measure the distance between a chosen initial condition and a slightly perturbed version of it as a function of time. If the distance in phase space grows exponentially in time, this is an indication of chaotic dynamics. In general, there is a spectrum of Lyapunov exponents, but the dynamics are dominated by the maximum Lyapunov exponent, which can be empirically measured. In the inset plots of FIG. 4a and FIG. 4b, we show maximum Lyapunov exponents for the 2D Ising model and the 3D spin glass problem, respectively. We observe that in both cases the log distance acquires a linear slope and then saturates. The linear and positive slope exhibits chaotic dynamics before saturating. The saturation of the distance is indicative of the finite size of the phase space where a distance between two trajectories has an upper bound.
The specifics of the calculation are as follows: the phases $\mathbf{x}$ and the states $\mathbf{m}$ are randomly initialized. We then introduce a perturbation in the initial conditions by offsetting the phase $x_{i}$ of a randomly selected c-bit by a magnitude of $0.1$. We then consider the Euclidean distance between the phases of the perturbed and unperturbed systems: $D=||\mathbf{x}-\mathbf{x}^{\prime}||$, where $\mathbf{x}^{\prime}$ represents the perturbed system’s phases. The log Euclidean distance is plotted over 200 time steps, with the system at constant inverse temperature $\beta=1$. The slope represents the Lyapunov exponent $\lambda$, as $D\sim\exp{\lambda t}$. The results are averaged over 100 trials, where each trial compares one unperturbed system $\mathbf{x}$ with 10 perturbed versions, $\mathbf{x}^{\prime}$.